# Isotope shift analysis with the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition in ytterbium

### v1

@@ IV.4 Search for new bosons
Together with theoretically calculated $D_{\alpha}$ and the linearity in the King plot for the 431 nm transition, a constraint on the existence of new bosons mediating the force between an electron and a neutron is set. It should be noted that with the relative uncertainty of $\sim 1\times 10^{-5}$ for the isotope shifts in the 431 nm transition, simple analysis with a two-dimensional King plot is good enough, as other sources of uncertainties such as the ones on $\mu^{AA^{\prime}}$ is orders of magnitude smaller. To be most conservative, the effect of the nuclear deformation along the blue line in Fig. 4 is first subtracted from the data point for the 431 nm transition, so that all nonlinearity would arise from the new bosons. As shown in Fig. 4, the effect of the new boson on the $\zeta_{+}-\zeta_{-}$ plane has a slope of -0.2585. On this line, $-6.685\times 10^{5}<v_{ne}D_{431,578}<2.182\times 10^{6}$ is the allowed region for the new bosons with 95% confidence level, which includes $v_{ne}=0$. Thus, such bosons are excluded in the region beyond this, as plotted in Fig. 5. The constraint is more than an order of magnitude worse compared to the most stringent constraint reported in Ref. [1].
With an improvement in the accuracy of the isotope shift measurement for the 431 nm transition, it is expected to increase the sensitivity to the region that has never been investigated before. Fig. 5 shows the sensitivity for the improved accuracy by three orders of magnitude (down to $\sim 18$ Hz), with an assumption that the uncertainty in the isotope shifts for the 431 nm transition is the primary source of the uncertainty. At this point, analysis with three transitions described in Ref. [6] is necessary.
Figure 5: Constraint on the existence of new bosons mediating the force between an electron and a neutron: the black region is excluded by electron $g-2$ measurements and neutron scattering experiments [3]. The red region is the most stringent exclusion reported so far in Ref. [1], obtained from the analysis with the 436, 467, and 578 nm transitions. The blue region is the excluded region by the 431 and 578 nm transitions obtained in this work. The blue dashed line is the expected sensitivity when the accuracy in isotope shifts for the 431 nm is improved by a factor of 1000.
@@ V Conclusion
The full list of absolute frequencies for the 431 nm transition in ${}^{170,172,173,174,176}$b is reported, and various analyses based on the data are performed. The King plot analysis did not show as much nonlinearity as reported for other transitions. On one hand, this can ensure the absence of any significant systematic shifts in the measurement, but on the other hand, further improvement in accuracy is desired to investigate the effects inducing the nonlinearity, including the new bosons mediating the force between an electron and a neutron. The hyperfine constants are precisely determined. However, the observed octupole moment is not statistically significant, and this also motivates more precise spectroscopy. The difference in root-mean-square nuclear charge radii between isotopes is determined by averaging isotope shifts of multiple transitions, supporting smaller values compared to some of the previous reports.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI JP21K20359, JP22H01161, JP22K04942, JST FOREST JPMJFR212S and JST-MIRAI JPMJM118A1. We are grateful to D. Akamatsu, K. Hosaka, H. Inaba, and S. Okubo for the development of the frequency comb and the stable laser at 1064 nm. A. K. acknowledges the partial support of a William M. and Jane D. Fairbank Postdoctoral Fellowship of Stanford University for the early stage of this project.
Appendix A Summary of the result of theoretical calculation
[…]
this work Ref. [1] Ref [22] Experiment
$\nu_{361}$ 875.19 819.47 829.76
$\nu_{399}$ 744.82 751.53
$\nu_{411}$ 691.88 707.00 729.48
$\nu_{431}$ 810.67 695.17
$\nu_{436}$ 649.86 679.86 688.35
$\nu_{467}$ 727.06 1051.44 642.12
$\nu_{556}$ 524.06 543.18 539.39
$\nu_{578}$ 500.53 522.78 522.68 518.29
$F_{361}$11.988 -13.528
$F_{399}$7.020
$F_{411}$15.534 -14.715
$F_{431}$ 17.771
$F_{436}$15.770 -14.968
$F_{467}$ 35.260 36.218
$F_{556}$9.893 -10.951(21)
$F_{578}$9.692 -9.719 -10.848(21)
$K_{361}$926
$K_{399}$ 412
$K_{411}$308 -752
$K_{431}$ 10172
$K_{436}$158 -661
$K_{467}$ 10167 12001
$K_{556}$532 -280(72)
$K_{578}$527 -288(75)
$D_{361}$ 5.6683
$D_{399}$
$D_{411}$ 43.37 43.158
$D_{431}$94.34
$D_{436}$ 49.74 48.634
$D_{467}$295.06 -352.38
$D_{556}$
$D_{578}$39.49 -42.855
$F_{399}/F_{578}$ 0.7243 0.4453(12)
$F_{431}/F_{578}$1.8336 -1.62301(8)
$F_{556}/F_{578}$ 1.0207 1.00594(26)
$K_{399,578}$ 793.7 1332(17)
$K_{431,578}$ 9205.7 6430.9(1.4)
$K_{556,578}$ 5.929 33.2(4.0)
$D_{431,578}$266.7
Appendix B Table for $\langle r^{2}\rangle^{AA^{\prime}}$
Values for $\langle r^{2}\rangle^{AA^{\prime}}$ that were used to generate Fig. 2 are summarized in Table 7.
Table 7: Summary of $\langle r^{2}\rangle^{AA^{\prime}}$. The second column is the valued obtained from the isotope shift of the 431 nm transition. The third column shows the average over the 399, 431, 556, and 578 transition data. The uncertainties for the second column are originated from the spectroscopy only, and the uncertainties in the third column are the standard deviation over the values derived from four transitions.
[…]
431 nm average
168 -0.474 791 6(21) -0.442(78)
170 -0.335 961 7(22) -0.312(54)
171 -0.286 988 3(25) -0.267(47)
172 -0.205 025 1(23) -0.189(33)
173 -0.157 682 6(23) -0.146(25)
174 -0.100 263 6(22) -0.093(16)
176  0  0
Appendix C Details of the $g$ factor determination
To determine the $g$ factors for the $F=7/2$ and $5/2$ states in ${}^{173}$b, the bias magnetic field $B_{\parallel}$ along the direction of the propagation of the 431 nm probe beam is varied to obtain the amount of Zeeman splitting. These are shown in Fig. 6. The plot is fitted with the fit function $\Delta\nu=g_{F}\mu_{B}\sqrt{(B_{\parallel}-B_{0})^{2}+B_{\perp}^{2}}$ to accomodate the transverse magnetic field $B_{\perp}$, where $g_{F}$, $B_{0}$, and $B_{\perp}$ are fitting parameters. $g_{F}$ is converted to $g_{J}$ with the relation
@@ Appendix C Details of the $g$ factor determination
(a) The $F=7/2$ state
(b) The $F=5/2$ state
(c) The $F=3/2$ state
Figure 6: The amount of Zeeman splitting $\Delta\nu$ for applied bias magnetic field $B_{\parallel}$ for hyperfine states in ${}^{173}$b. The uncertainties on the final values for $g_{F}$ in the main text are inflated by $\sqrt{\chi^{2}/{\rm ndf}}$. if $\sqrt{\chi^{2}/{\rm ndf}}>1$.
Appendix D Details of the absolute frequency determination
Several spectra are taken for the absolute frequency measurement for each isotope and hyperfine state to confirm the consistency of the data. Weighted average of these data are regarded as the absolute frequencies, as shown in Fig. 7. The statistical uncertainty is inflated by $\sqrt{\chi^{2}/{\rm ndf}}$ if $\sqrt{\chi^{2}/{\rm ndf}}>1$, and then merged with the systematic uncertainties.
(a) ${}^{176}$b
(b) ${}^{174}$b
(c) ${}^{172}$b
(d) ${}^{170}$b
(e) the $F=9/2$ state in ${}^{173}$b
(f) the $F=7/2$ state in ${}^{173}$b
(g) the $F=5/2$ state in ${}^{173}$b
(h) the $F=3/2$ state in ${}^{173}$b
(i) the $F=1/2$ state in ${}^{173}$b
Figure 7: Plots for determining absolute frequencies: $\nu_{0}$ in the box shows the weighted average. The uncertainty on $\nu_{0}$ is without the inflation by $\sqrt{\chi^{2}/{\rm ndf}}$, whereas the green band is with the inflation if $\sqrt{\chi^{2}/{\rm ndf}}>1$.
Figure 8: Plots of the quadrupole moment $Q$ in Table 4, hyperfine anomaly $\Delta_{171,173}/10^{-3}$, hyperfine constants $A_{h}^{171}$, $B_{h}^{173}$, $B_{h}^{173}/A_{h}^{173}$, and field shift $F_{\alpha}$ against $\delta\nu^{171}$: three points correspond to the 399, 431, and 556 nm data. The red line is the linear fit.
Figure 9: Plots of the quadrupole moment $Q$ in Table 4, hyperfine anomaly $\Delta_{171,173}/10^{-3}$, hyperfine constants $A_{h}^{171}$, $B_{h}^{173}$, $B_{h}^{173}/A_{h}^{173}$, and field shift $F_{\alpha}$ against $\delta\nu^{171}$: three points correspond to the 399, 431, and 556 nm data. The red line is the linear fit.
Figure 10: Plots of the quadrupole moment $Q$ in Table 4, hyperfine anomaly $\Delta_{171,173}/10^{-3}$, hyperfine constants $A_{h}^{171}$, $B_{h}^{173}$, $B_{h}^{173}/A_{h}^{173}$, and field shift $F_{\alpha}$ against $\delta\nu^{173}/\delta\nu^{171}$: three points correspond to the 399, 431, and 556 nm data. The red line is the linear fit.
Appendix E Details of the analysis for the residual from the King plot
Plots of $Q$, $\Delta_{171,173}/10^{-3}$, $A_{h}^{171}$, $B_{h}^{173}$ , $B_{h}^{173}/A_{h}^{173}$, and $F_{\alpha}$ against $\delta\nu^{171}$, $\delta\nu^{173}$, and $\delta\nu^{173}/\delta\nu^{171}$ are shown in Figs. 8, 9, and 10.
Appendix F Formula for hyperfine structure
The absolute frequency for each hyperfine level in ${}^{173}$b satisfies the following equation.
$\nu$ $=$ $\nu_{\rm ave}+A_{h}^{173}\frac{K}{2}+B_{h}^{173}\frac{3K(K+1)4I(I+1)J(J+1)}{8I(2I-1)(2J-1)}$ (11)
@@ Appendix F Formula for hyperfine structure
with $\nu_{\rm ave}$, $A_{h}^{173}$, $B_{h}^{173}$, and $C_{h}^{173}$ being constants that can be determined by measurements, $I$ being nuclear spin, $J$ being total electronic angular momentum, and $K=F(F+1)-I(I+1)-J(J+1)$, with ${\bf F}={\bf I}+{\bf J}$. The 2nd, 3rd, and 4th terms correspond to the magnetic dipole, electric quadrupole, and magnetic octupole components, respectively.

### v2

@@ IV.4 Search for new bosons
Together with theoretically calculated $D_{\alpha}$ and the linearity in the King plot for the 431-nm transition, a constraint on the existence of new bosons mediating the force between an electron and a neutron is estimated. It should be noted that with the relative uncertainty of $\sim 1\times 10^{-5}$ for the isotope shifts in the 431-nm transition, simple analysis with a two-dimensional King plot is good enough, as other sources of uncertainties such as the ones on $\mu^{AA^{\prime}}$ are orders of magnitude smaller. To be most conservative, the effect of the nuclear deformation along the blue line in Fig. 4 is first subtracted from the data point for the 431-nm transition, so that all nonlinearity would arise from the new bosons. As shown in Fig. 4, the effect of the new boson on the $\zeta_{+}-\zeta_{-}$ plane has a slope of -0.2585. On this line, $-6.685\times 10^{5}<v_{ne}D_{431,578}<2.182\times 10^{6}$ is the allowed region for the new bosons with 95% confidence level, which includes $v_{ne}=0$. Thus, such bosons are excluded in the region beyond this, as plotted in Fig. 5. The constraint is more than an order of magnitude worse compared to the most stringent constraint reported in Ref. [3].
An improved accuracy of the isotope-shift measurement for the 431-nm transition increases the sensitivity to the region that has never been investigated before. Fig. 5 shows the sensitivity for the improved accuracy by three orders of magnitude (down to $\sim 18$ Hz), with an assumption that the uncertainty in the isotope shifts for the 431-nm transition is the primary source of the uncertainty. At this level, analysis with three transitions described in Ref. [8] is necessary.
Figure 5: Constraint on the existence of new bosons mediating the force between an electron and a neutron: the black region is excluded by electron $g-2$ measurements and neutron-scattering experiments [5]. The red region is the most stringent exclusion reported so far in Ref. [3], obtained from the analysis with the 436-, 467-, and 578-nm transitions. The blue region is the excluded region by the 431- and 578-nm transitions obtained in this work. The blue dashed line is the expected sensitivity when the accuracy in isotope shifts for 431 nm is improved by a factor of 1000.
@@ V Conclusion
The full list of absolute frequencies for the 431-nm transition in 170,172,173,174,176Yb is reported, and various analyses based on the data are performed. The King plot analysis did not show as much nonlinearity as reported for other transitions. On one hand, this can ensure the absence of any significant systematic shifts in the measurement, but on the other hand, further improvement in accuracy is desired to investigate the effects inducing the nonlinearity, including the new bosons mediating the force between an electron and a neutron. The hyperfine constants are precisely determined. However, the observed octupole moment is not statistically significant, and this also motivates more precise spectroscopy. The difference in root-mean-square nuclear charge radii between isotopes determined by our analysis supports smaller values compared to some of the previous reports.
@@ Acknowledgements.
This work was supported by Japan Society for the Promotion of Science KAKENHI Grants No. JP21K20359, JP22H01161, JP22K04942, and Japan Science and Technology Agency FOREST Grant No. JPMJFR212S and JST-MIRAI Grant No. JPMJM118A1. We are grateful to D. Akamatsu, K. Hosaka, H. Inaba, and S. Okubo for the development of the frequency comb and the stable laser at 1064 nm. A. K. acknowledges the partial support of a William M. and Jane D. Fairbank Postdoctoral Fellowship of Stanford University for the early stage of this project.
Figure 6: The amount of Zeeman splitting $\Delta\nu$ for applied bias magnetic field $B_{\parallel}$ for the (a) $F=7/2$, (b) $F=5/2$, and (c) $F=3/2$ hyperfine states in 173Yb. The uncertainties on the final values for $g_{F}$ in the main text are inflated by $\sqrt{\chi^{2}/{\rm ndf}}$. if $\sqrt{\chi^{2}/{\rm ndf}}>1$.
Appendix A Details of the $g$ factor determination
To determine the $g$ factors for the $F=7/2$, $5/2$, and $3/2$ states in 173Yb, the bias magnetic field $B_{\parallel}$ along the direction of the propagation of the 431-nm probe beam is varied to obtain the amount of Zeeman splitting. These are shown in Fig. 6. The plot is fitted with the fit function $\Delta\nu=g_{F}\mu_{B}\sqrt{(B_{\parallel}-B_{0})^{2}+B_{\perp}^{2}}$ to accommodate the transverse magnetic field $B_{\perp}$, where $g_{F}$, $B_{0}$, and $B_{\perp}$ are fitting parameters. $g_{F}$ is converted to $g_{J}$ with the relation
@@ Appendix A Details of the $g$ factor determination
Appendix B Formula for hyperfine structure
To the extent of the first-order interaction, the absolute frequency $\nu_{F}$ for the hyperfine level with total angular momentum $F$ in 173Yb satisfies the following equation.
$\nu_{F}$ $=$ $\nu_{\rm ave}+A_{h}^{173}\frac{K}{2}+B_{h}^{173}\frac{3K(K+1)-4I(I+1)J(J+1)}{8I(2I-1)J(2J-1)}$ (8)
@@ Appendix B Formula for hyperfine structure
with $\nu_{\rm ave}$, $A_{h}^{173}$, $B_{h}^{173}$, and $C_{h}^{173}$ being constants that can be determined by measurements, $I$ being nuclear spin, $J$ being total electronic angular momentum, $K=F(F+1)-I(I+1)-J(J+1)$, and ${\bf F}={\bf I}+{\bf J}$. The second, third, and fourth terms correspond to the magnetic dipole, electric quadrupole, and magnetic octupole components, respectively.
The formalism for the second-order hyperfine interaction is described in Ref. [49]. Briefly, the shift is described as
$W_{F}^{(2)}=\sum_{\gamma^{\prime}J^{\prime}}\frac{|\langle\gamma^{\prime}I^{\prime}J^{\prime}FM_{F}|H_{HFI}|\gamma IJFM_{F}\rangle|}{E_{\gamma J}-E_{\gamma^{\prime}J^{\prime}}},$ (9)
where $\gamma$ is the set of quantum numbers for electronic orbitals other than $I$, $J$, $F$, and $M_{F}$; $H_{\rm HFI}$ is the Hamiltonian for the hyperfine interaction; and $E_{\gamma J}$ is the energy for the state $\gamma J$. When the case of $J=2$ is considered, contributions from the adjacent fine structures dominate the sum over $\gamma^{\prime}J^{\prime}$. When we consider only the dipole-dipole interaction and dipole-quadrupole interactions, Eq. (9) results in
$W_{F}^{(2)}=$ |$\left|\begin{Bmatrix}F&J&I\\ 1&I&J-1\end{Bmatrix}\right|^{2}\eta$
+$+\begin{Bmatrix}F&J&I\\ 1&I&J-1\end{Bmatrix}\times\begin{Bmatrix}F&J&I\\ 2&I&J-1\end{Bmatrix}\zeta.$
$\eta$ and $\zeta$ are $F$-independent terms defined as
$\eta=$ $\frac{(I+1)(2I+1)}{I}\mu^{2}\frac{|\langle\gamma J-1||T_{1}^{e}||\gamma J\rangle|^{2}}{E_{\gamma J}-E_{\gamma J-1}}$
$\zeta=$ $\frac{(I+1)(2I+1)}{I}\sqrt{\frac{2I+3}{2I-1}}$
×$\times\mu Q\frac{\langle\gamma J-1||T_{1}^{e}||\gamma J\rangle\langle\gamma J-1||T_{2}^{e}||\gamma J\rangle}{E_{\gamma J}-E_{\gamma J-1}}$
with $T_{1}^{e}$ the spherical tensor of rank 1 acting on electronic coordinates. This leads to
$W_{F}^{(2)}=\left\{\begin{array}[]{ll}0&(F=9/2)\\ \frac{3}{140}\eta+\frac{3}{140\sqrt{2}}\zeta&(F=7/2)\\ \frac{32}{1575}\eta-\frac{8}{525\sqrt{2}}\zeta&(F=5/2)\\ \frac{1}{100}\eta-\frac{1}{50\sqrt{2}}\zeta&(F=3/2)\\ 0&(F=1/2)\end{array}\right.$ (15)
The contribution of the second-order hyperfine interaction to the A, B, C, and D constants can be calculated by first writing down $\nu_{F-1}-\nu_{F}$ ($F=3/2,5/2,7/2,9/2$) as a function of hyperfine constants. Solving this for the hyperfine constants, the hyperfine constants are described as a function of $\nu_{F-1}-\nu_{F}$. Using the same inverse matrix on $W_{F-1}^{(2)}-W_{F}^{(2)}$, the following contributions of the second-order hyperfine interaction on the A, B, and C constants are obtained:
$A:$ $\frac{1}{1050}\eta-\frac{2\sqrt{2}}{2625}\zeta$
$B:$ $\frac{8}{315}\eta+\frac{\sqrt{2}}{105}\zeta$
$C:$ $\frac{1}{350\sqrt{2}}\zeta$
In the analysis in the main text, $W_{F}^{(2)}$ described in Eq. (15) is added to the right-hand side of of Eq. (8).